XxLinaxX:

A bakery sells muffins for $3.50 each. A beverage is $1.75. A class purchases 32 items and spends a total of $87.50. Define your variables. Write the system of equations and represent it as a matrix equation. State the value of the determinant. Use matrices to solve the system. Find the number of muffins and the number of beverages purchased.

dude:

The prices mentioned in the question connect to the variables Use a variables for muffins and beverages You were asked to make 2 equations. A hint here is to create one for price and another for amount of items bought

XxLinaxX:

...no comment

XxLinaxX:

tbh that doesn't help my slow self.

dude:

Do you know what variables are?

XxLinaxX:

yes, they are x y z etc.

XxLinaxX:

so x for muffins and y for beverages?

dude:

I meant in the question but yes youre right

XxLinaxX:

smh-

Mercury:

if you let x = # of muffins and y = # of beverages the amount of $$ spent on muffins is (price of muffins)(number of muffins) same logic w/ the beverages. (price of beverages)(number of beverages) adding the amt spent on beverages + muffins will give you the total cost plugging in the appropriate values, 3.5x + 1.75y = 87.5

Mercury:

the class purchases 32 items, so x + y = 32 your system: 3.5x + 1.75y = 87.5 x + y = 32 now to convert this to matrix form simply write a 2x2 matrix with the coefficients 3.5, 1.75, 1, and 1 in that order multiply that by a 2x1 matrix with items x and y write an equals sign then another 2x1 matrix with the constants 87.5 and 32

Mercury:

the determinant of the 2x2 coefficient matrix is simply 

Mercury:

solving the system: first find the inverse of the coefficient matrix A  where |A| is the determinant from the previous step multiply both sides of the system by the inverse matrix ***important*** the inverse matrix must be written on the left on both sides, a.k.a A^(-1) * A [x,y] = A^(-1) [constant matrix] the inverse cancels out with A and you just have [x,y] = solution matrix